OpenStudy (yttrium):

The range of a projectile thrown horizontally is said to be dependent to the following quantities 'R'~f(v. H, g, m). Derive an expression for R in terms of v, H, g and m. (I got R = V^2/g but I know if I am just correct because I don't really know the formula for the range of projectiles, can somebody verify?) Thanks!

OpenStudy (yttrium):

@hartnn , can you help me?

OpenStudy (john_es):

The solution should depends on H. Do you put in your kinematic equations the initial height as H?

OpenStudy (yttrium):

Yes. But it appear like this in the end. Btw, can you start it? So I can verify if I was in a right path.

OpenStudy (anonymous):

for x: x = vt so we have to find the flight time t. for y: 0 = h - 0.5gt^2 now you can find t using y equation and plug it into x equation and have R.

OpenStudy (anonymous):

indeed x = vsqrt(2h/g) using dimensional analysis you can assume that the range will depend on v g and h but not on m since nothing will cancel the mass units.

OpenStudy (john_es):

Exact.

OpenStudy (yttrium):

Is the solution you made is dimensional analysis approach? @Coolsector

OpenStudy (anonymous):

no

OpenStudy (yttrium):

But how can we do it as dimensional analysis approach?

OpenStudy (anonymous):

using dimensional analysis we would not know that there is a "2" in the square root

OpenStudy (yttrium):

Can you show me how to start?

OpenStudy (anonymous):

well i think knowing the answer make me cheat here : units of distance [m] , time [s] . we want g, h, and v in the expression: x ~ v^(a) * g^(b) * h^(c) = [m/s]^a * [m/s^2]^b * [m]^c = [m] so a+2b = 0 b = -a/2 and a+b+c = 1 c = 1-a/2 maybe i miss something ( im not really use to dimensional analysis) but i think we should make another assumption about a (the power of v) if we take it to be 1 we get the correct powers

OpenStudy (anonymous):

assuming a =1 though is reasonable due to x=vt in such throw

OpenStudy (yttrium):

Thanks for telling we can assume a=1 In my solution, I did R ~ v^w h^x g^y m^z [R] = L [v] = LT^-1 [H] = L [g] = LT^-2 [m] = M Therefore, K= (LT^-1)^W L^x (LT^-2)^y M^z L = L^(w+x+y)T^(-w-2y)(M^z) z=0 w = -2y w+x+y = 1 Letting w = 1, I get y = -1/2 x = 1/2 Therefore, R ~ VH^1/2g^-1/2 or \[R ~ \frac{ v \sqrt{H} }{ \sqrt{g} }\] which is \[R = V \sqrt{\frac{ H }{ g }}\] HAHAHA! I got it. Thanks, but how letting w= 1 reasonable??

OpenStudy (anonymous):

as i said: "assuming a =1 though is reasonable due to x=vt in such throw" i hope it is correct and not "cheating" anyway you cant have an equal sign .. using dimensions you can only have proportional..

OpenStudy (yttrium):

Okay. I think it's true anyway. Thanks! :)

OpenStudy (anonymous):

about the proportional : as you can see we didnt get the 2 under the square root as we got in the "x = vsqrt(2h/g)" which is the correct full answer

OpenStudy (yttrium):

@luigi0210 @nincompoop can I really substitute 1 in the w in my solution here?

OpenStudy (yttrium):

@psymon